if y varies directly as x and inversely as z, find y whe x=3 and z=2, if y=90 when x=30 and z=1.5
Answer by scott8148(6628)

y * z = k * x

90 * 1.5 = k * 30 ___ k = 4.5

y * 2 = 4.5 * 3 ___ y = 6.75